OpenStudy (anonymous):

Which of the following points are NOT on the graph of y = -2x + 3? a) (1, 1) b) (-1, 5) c) (-2, 3) d) (2, -1)

OpenStudy (kidrah69):

\(\huge\color{orange}{WELCOME~TO~OPENSTUDY!!~:)}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an image tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/

OpenStudy (kidrah69):

All you do is plug in each coordinate if the numebers are a solution then it will be on the graph if not that is your answer :)

OpenStudy (kidrah69):

1) \[1= -2(1) + 3\] 2) \[5= -2(-1) + 3\] 3) \[3 = -2(-2)+ 3\] 4) \[-1 = -2(2)+ 3\]

OpenStudy (kidrah69):

Now all you have to do is solve =)